OpenStudy (anonymous):

Jeannette's starting salary is $25,000 a year. She will get a 4% raise annually. How much money will she have made after 7 years? A. $165,824.39 B. $197,457.36 C. $230,355.66 D. $256,355.66

OpenStudy (kropot72):

Use the formula for a geometric progression: \[S _{n}=\frac{a(r ^{n}-1)}{r-1}=\frac{25000(1.04^{7}-1)}{1.04-1}=you\ can\ calculate\]